How many grams of a 25% antibiotic solution should be added to 10 grams of a 60% antibiotic solution in order to get a 30% antibiotic solution?
Answer by checkley77(12844)

.60*10+.25x=.30(10+x)
6+.25x=3+.3x
.25x-.30x=3-6
-.05x=-3
x=-3/-.05
x=60 grams of 25% solution are used.
Proof:
.60*10+.25*60=.30(10+60)
6+15=.30*70
21=21